You want to earn $60 per month on a 9.5% investment. how much do you need to invest? Please show steps
Answer #1

Since time is not given, i am assuming this is a perpetuity.

Rate of return = 0.095 / 12 = 0.007917 or 0.7917%

Present value = Cash flow / rate of return

Present value = 60 / 0.007917

Present value = $7,578.63

Need to invest $7,578.63

If you have more information like number of periods for the cash flow, please post it in comments. I will redo the question.